# Revision of two-temperature magnetically arrested flows onto a black hole

### v1

@@ 2.4 Effects of grid resolution
Table 1 lists four test two-temperature MAD $a_{*}=0$ models carried out at different numerical resolutions. Figure 1 shows equatorial cuts through the four models at final integration time, $t_{f}=10,000$M. Figure 2 displays the corresponding light-curves for Stokes ${\mathcal{I}}$, RM, LP and CP calculated for time interval between 5, [^0] M and 10, [^0] M at three observing frequencies assuming K model for electron temperatures. In Figure 2 bottom panels display a comparison of $M_{3}$, RM, LP, CP distributions at 229 GHz as a function of grid resolution. There is a very good agreement between $M_{3}$, RM, LP and CP distributions. Although not shown, this is also true for two other neighboring frequencies of 86 and 690 GHz at which the modeled source could be observed.
We conclude that the results presented in the remaining part of this work do not strongly depend on the chosen GRMHD grid resolution. Around the current level of resolutions, which are typically used for building EHT GRMHD model libraries, all modeled quantities are rather settled and the small variations in observables are due to different realizations of turbulence in individual runs [^1] Turbulence is always initialized by a random perturbation of the gas internal energy.. We note that we cannot exclude that our models are located at the resolution plateau and that increasing resolution 10 times in each direction may result in the simulation better resolving more physical effects (e.g., better resolution of shocks and reconnection layers may lead to different electron heating in different regions of the flow) that could lead to quantitative and qualitative changes in the emission properties.
@@ 3 Results
3.1 Turbulent heating vs. R($\beta$) model
Parameters of all the fiducial runs are summarized in Table 1. Figure 3 shows an example of snapshots from our fiducial two-temperature MAD $a_{*}=0.9375$ simulation. The fiducial light-curves produced by the GRMHD simulations with different parameters are presented in Appendix A and here we discuss distributions of quantities calculated based on these light-curves.
Figure 4 displays correlations of observables in models R1-160 and K for the default viewing angle $i=160$ deg, combined for black hole spins $a_{*}=0,0.5,0.9375$. In $M_{3}$ panels, K models are moderately correlated with R1-160 models and typically K models are less variable compared to models R1-160. Spectral slope $\alpha_{I}$ and CP in model K are highly correlated with those of model R10. This is not the case for RM and for LP for which correlation is much weaker.
Figure 5 shows comparison of distributions of observables for models shown in Figure 4. $M_{3}$, LP and CP averages (and standard deviations) in models K are weakly distinguishable from those in models R1-160.
It is useful to look at properties of maps used to synthesize the integrated observables. Figure 6 shows examples of EHT-like images at 229 GHz produced by a random snapshot of $a_{*}=0.5$ simulation with R1,R10,R160 and K electron models for the fiducial viewing angle of $i=160$ deg. Following Event Horizon Telescope Collaboration et al. 2024 we characterize these images using the following metrics. Net linear and circular polarizations are defined as:
@@
The image-averaged linear and circular polarizations defined as:
@@
where sums are carried out over all image pixels. Geometry of linear polarization is described with amplitudes and phases of a complex function
@@
is the complex polarization vector and where $\rho,\phi$ are the polar coordinates in the image plane (see Palumbo et al. 2020 or Event Horizon Telescope Collaboration et al. 2024 for details). All these metrics are displayed in the example snapshot in Figure 6. Here it is more evident that the resolved images of models K are most similar to models R10 in both LP and CP. Models R1 differ from K in CP maps. Models R40-160 have already distinct appearance compared to model K, both in LP and CP, as also evident from the comparison of integrated CP. When scored against the existing or a future (total intensity or polarimetric) EHT data, models K may favor similar accretion flow and geometrical parameters as models R10.
It is important to show that the $T_{\rm p}/T_{\rm e}$ in K models does not precisely follow the $R(\beta)$ model R10, regardless of $Q^{+}_{\rm p}/Q^{+}_{\rm e}$ being strongly dependent on $\beta$. Example of $T_{\rm p}/T_{\rm e}$ maps for a single snapshot of MAD model with high spin is shown in Figure 7 (top and middle panels). The discrepancy between $T_{\rm p}/T_{\rm e}$ in model R10 and K indicates that there is no simple relationship between $Q^{+}_{\rm p}/Q^{+}_{\rm e}$ and $T_{\rm p}/T_{\rm e}$ as already shown by Ressler et al. (2015) for non-MAD models. Hence, the exact electron temperature consistency is not the main reason behind the similarity of K and R10 models. The difference in $T_{\rm p}/T_{\rm e}$ between the models is the smallest close to the black hole and it increases with radius. The emitting regions close to the black hole have on average $T_{\rm p}/T_{\rm e}$ =10. R10 and K models therefore share the scaling ${\mathcal{M}}$ unit to reproduce the same flux meaning that $\rho$ and $B$ in the synchrotron radiative transfer coefficients are identical leading to similar images. The differences between the model’s emissions reflect differences in $\Theta_{e}$ (shown in lower panels in Figure 7) and they seem small from the observational point of view. Due to the Faraday rotation effects LP and RM are the most sensitive among our observables to electron temperatures at small and larger radii. This explains the weaker correlation of LP and RM between R10 and K models visible in Figure 4.
@@ 3.2.1 Effects of radiative cooling
We carry out radiative GRMHD (GRRMHD) simulation of MAD $a_{*}=0$ model to check impact of radiative cooling on electron temperatures and emission properties. The radiative processes included in GRRMHD simulation include synchrotron emission, self-absorption, Compton scatterings and bremsstrahlung. To reduce the large computational cost of models with activated multi-frequency radiative transfer the interactions between photon field and gas is allowed only within $100$M. Any radiative transfer effects outside of this sphere are negligible. When producing emission the code takes into account the fact that the combined four-momentum of gas and radiation has to be conserved. Emission/absorption/scattering can therefore exert force on gas and change the electron thermodynamics (radiative cooling or heating). In the models considered here, the average optical thickness of gas is $\tau\ll 1$ and luminosities are strongly sub-Eddington, $L\ll L_{\rm Edd}$, hence it is expected that in models with activated radiative transfer mainly electron thermodynamics is affected and mainly due to synchrotron cooling. Other effects are weak for the considered accretion rates. In Appendix B we present quality factors for the GRRMHD simulation to demonstrate that the interactions between gas and radiation are resolved by adequate number of Monte Carlo samples.
@@
Figure 8 displays volume averaged radial profiles of density, magnetic field, $T_{\rm p}/T_{\rm e}$ (or R), and $\Theta_{e}$ in GRMHD and GRRMHD model with K heating mechanism and ${\mathcal{M}}=5\times 10^{11}$ at different time moments, earlier and later in the evolution. All quantities are additionally weighed by $\rho B$ factor to show quantities averaged over regions where most of the synchrotron emission comes from. There is a small difference in $\Theta_{e}$ and $R$ radial profiles in models with and without radiative effects, and the difference is due to different realization of turbulence in each model. Unsurprisingly, at later times in both models $R\sim 10$ for $r<20$M, consistent with results in Section 3.1. Figure 9 additionally shows that there is no difference in radiative characteristics between GRMHD (K) and GRRMHD (Kcool) models based on $15,000-30,000$M lightcurves.
@@
Since the test model with cooling has the highest ${\mathcal{M}}$ among all K models considered in this work, all other K models shown in the paper are expected to cool down even slower than the test case. Radiative cooling effects in MADs with turbulent heating may be neglected when modeling Sgr A* millimeter emission.
3.2.2 Effects of non-thermal electrons
@@
Figure 10 shows comparison of observational characteristics of thermal and non-thermal K fiducial runs. Non-thermal models are slightly less variable compared to the thermal models. Spectral slope $\alpha_{I}$, RM, and LP of both thermal and non-thermal models are strongly correlated (the correlation is much stronger in comparison to $R(\beta)$ models). It is evident that the total intensity and linear polarization are not sensitive to details of the distribution function for the considered part of electromagnetic spectrum. CP, on the other hand, is different in both models, hence it may be used to discriminate between distribution functions. We note that CP is usually a more permissive observational constraint compared to LP (see e.g., Event Horizon Telescope Collaboration et al. 2024). Overall, with exception for CP, all radiative properties of our simple non-thermal models are tightly correlated with their purely thermal counterparts.
@@
Having checked how the results depend on numerical resolution effects and two most relevant physical effects, we can now compare the thermal two-temperature models directly to observations of plasma around Sgr A*.
Figure 11 shows comparison of prograde models K, at various viewing angles, to Sgr A* observables detected by ALMA at 86 and 213-229 GHz in April 2017 (Wielgus et al., 2022a, b, 2024). The modulation index at both frequencies increases with black hole spin and with inclination angle. In general all models are too variable to match the observed variability (this is also the case for R1-160 models as found in Event Horizon Telescope Collaboration et al. 2022 and shown here in Figure 5) but data strongly favor lower spins at lower inclination angles for which $M_{3}$ is the smallest. All models are too optically thin when compared to observations (as is also the case for R1-160 models, compare our Figure 5 with Figure 12, see also Ricarte et al. 2023). All models also require external Faraday screen to explain the observations of RM but we notice that models with low viewing angles ($i=160$ deg, $150$ deg) best recover the RM standard deviation towards the source. Linear polarization of $i=160-150$ deg models is consistent with both 229 and 86 GHz measurements. Models with $a_{*}=0$ recover CP at 229 GHz, but fail to recover CP at 86 GHz. On the other hand, models with $a_{*}=0.9375$ fail to recover CP at 229 GHz, but do recover CP at 86 GHz. Only model with $a_{*}=0.5$ does reasonably well with regards to CP at both frequencies. It is possible that a model with black hole spin in-between 0 and 0.5 would be a better fit to ALMA CP data. Among prograde MADs, model with $a_{*}=0.5$ is the closest to the ALMA observations altogether.
Although turbulence in the retrograde models reached quasi-stationary state within the emission zone, they are reported separately because they are integrated over shorter times compared to the prograde models. Figure 12 compares statistical properties of light-curves produced by the retrograde MADs. Two-temperature MAD with $a_{*}=-0.5$ performs better compared to MAD $a_{*}=-0.9375$ as they are less variable, but worse than MAD $a_{*}=0.5$, because prograde models better recover CP at both frequencies.
@@ 4 Discussions
In this work we have carried out several two-temperature MAD simulations with the main purpose of investigating impact of various numerical parameters on emission properties of these models, in particular grid resolution, but we also investigate impact of physical effects.
@@
As for physical parameters, radiative cooling can be safely omitted in Sgr A* two-temperature MAD models with turbulent heating, as radiative effects do not significantly affect the electron temperature evolution. These effects may be important when modeling M87, which is a brighter source. We have also carried out first exploratory survey of models in which turbulent heating is dissipated into a non-thermal electron distribution function. Our non-thermal models display very similar characteristics compared to the thermal models. However including physics of non-thermal electrons in the emission models slightly decreases the model variability and alters CP compared to purely thermal models. Other, also spatially varying, parameters of the $\kappa$ distribution function could be considered and change our result (e.g., Davelaar et al., 2019).
@@
Long-duration two-temperature simulations indicate that K thermal models are only roughly approximated with R10 models. Using total intensity observables Mizuno et al. (2021) found that the best match to the turbulent heating models are R5 models. Given that their electron temperatures are defined differently (compare their Eq. 8 with our Eq. 4) and are by a factor of $\sim 2$ smaller than ours, the best match with R10 found here is consistent with the previous finding.
@@
In Section 3.3 we show that none of our long-duration two-temperature K simulations recover general properties of light-curves of Sgr A* observed by ALMA ideally but the ALMA data favor models with lower viewing angles $i=150-160$ deg for which the variabilities of Stokes ${\mathcal{I}}$ and RM are the smallest. All models have major difficulty with recovering the observed model spectral index, in excellent agreement with previous two-temperature study by Dexter et al. 2020a. All models recover the observed net LP but require introduction of an additional external Faraday screen to recover the observed RM. Only $a_{*}=0.5$ model recovers Sgr A* net CP at two observed frequencies, simultaneously, and it is possible that a spin in between 0-0.5 would produce even better fitting results. The preferred moderate/low spin value is consistent with spin estimates based on two-temperature models of Dexter et al. 2020a, although for different than turbulent heating scenario, different initial conditions and different observational datasets. Here we also study evolution of two-temperature MADs for low and high spins for significantly longer time compared to Dexter et al. 2020a.
@@
It is possible that ALMA observes varying compact emission together with less varying extended emission that is not captured in our simulations (C.F. Gammie, private communication) hence it is preferable to compare models with EHT images (or best with the future EHT movies) rather than integrated quantities. When compared to EHT data, K models are expected to favor similar parameter space as R10 models. Polarimetric EHT data of Sgr A* have not yet ruled out all MAD R10 models. In fact 10 out of 20 passing MAD models are R10 models (see Fig. 8 in Event Horizon Telescope Collaboration et al. 2024). Based on similarity between R10 and K appearance, we argue that the turbulent electron heating scenario could be still consistent with the observations of Sgr A*. It is interesting to note that in Event Horizon Telescope Collaboration et al. 2024, MAD models R10 are favored for a set of intermediate viewing angles ($i=150-110$ deg) and positive and negative spin values. K models could narrow down the range of these two geometrical parameters.
@@
Based on joint total intensity and polarimetric EHT data, the current best-bet MAD model for Sgr A* is R160, not R10 (Event Horizon Telescope Collaboration et al., 2024). We ought to discuss the present results in context of other models for $Q^{+}_{\rm p}/Q^{+}_{\rm e}$. As shown by Dexter et al. (2020a) assuming Howes (2010) turbulent heating model for $Q^{+}_{\rm p}/Q^{+}_{\rm e}$, makes a negligible difference in the MAD electron temperatures compared to Kawazura et al. (2019) model. The reconnection $Q^{+}_{\rm p}/Q^{+}_{\rm e}$ model of Rowan et al. 2017, on the other hand, may result in slightly lower (but also around R10) electron temperatures with weaker plasma-$\beta$ dependency (Mizuno et al., 2021; Dexter et al., 2020a). Lower electron temperatures could make the MAD models optically thicker and less varying (Chan et al., 2024) hence pushing them towards the right direction, but this remains to be thoroughly examined. Weaker dependency on $\beta$ in models with reconnection heating may result in more disk-like emission of MADs rather than the disk/jet-like emission (Chael et al., 2018) making it more difficult to explain Sgr A* nearly flat radio spectrum (Mościbrodzka & Falcke, 2013). Radio spectral slope dependency on $Q^{+}_{\rm p}/Q^{+}_{\rm e}$ should be carefully checked with high (jet-resolving) resolutions two-temperature MAD models. Nevertheless, whichever $Q^{+}_{\rm p}/Q^{+}_{\rm e}$ scenario above (or a combination of scenarios) is selected, the two-temperature MAD models will have difficulty to naturally develop R160-like electron temperatures near the event horizon. Assuming that MAD is present in Sgr A* system, this may suggest that $R(\beta)$ models are inappropriate for MADs (we further briefly discuss prospects on how to resolve this issue in Appendix C). Alternatively, comparisons of time-varying models to time-averaged EHT images so far (Event Horizon Telescope Collaboration et al., 2022, 2024) introduce a strong bias in the electron temperature parameter inference.
@@
Our simulations are run with constant adiabatic index $\gamma_{\rm ad}=13/9$, which is not a self-consistent assumption in regions where electrons are mildly- or sub-relativistic. Recent simulations of Narayan et al. (2022) suggest that MAD dynamics is not affected by the choice of $\gamma_{\rm ad}$. It is unclear how different, or varying, $\gamma_{\rm ad}$ alters the discussed total intensity and polarimetric radiative properties of two-temperature MADs. This topic should be investigated in the future.
@@
Ideal GRMHD models considered here do not include many other, potentially important effects. Additional physics such as resistivity (e.g., Vos et al., 2024), viscosity, plasma composition (Emami et al., 2021; Wong & Gammie, 2022), anisotropic pressures (e.g., Foucart et al., 2017) or anisotropic electron distribution functions (e.g., Galishnikova et al., 2023), may all impact the radiative characteristics of MADs. Studying influence of all these effects on observed quantities goes beyond the scope of the current work. The current simulation set constitutes a reference point for future investigations.
@@
We thank Maciek Wielgus for providing ALMA data and comments on the paper. We also thank Andrew Chael, Charles Gammie, and Ben Ryan for many of their comments. We gratefully acknowledge the HPC RIVR consortium (www.hpc-rivr.si) and EuroHPC JU (eurohpc-ju.europa.eu) for funding this research by providing computing resources of the HPC system Vega at the Institute of Information Science (www.izum.si). Software ebhlight (Ryan et al., 2015, 2017, 2018), ipole (Mościbrodzka & Gammie, 2018), python (Oliphant, 2007), matplotlib (Hunter, 2007).

### v3

@@ 2.4 Effects of grid resolution
Table 1 lists four test two-temperature MAD models $a_{*}=0$ carried out at different numerical resolutions. Figure 1 shows the equatorial cuts through the four models at the final integration time, $t_{f}=10,000$M. Figure 2 displays the corresponding lightcurves for Stokes ${\mathcal{I}}$, RM, LP, and CP calculated for the time interval between 5000M and 10, [^0] M at three observing frequencies assuming the K model for electron temperatures. In Figure 2, the bottom panels display a comparison of the $M_{3}$, RM, LP, CP distributions at 229 GHz as a function of grid resolution. There is a very good agreement between the $M_{3}$, RM, LP and CP distributions. Although not shown, this is also true for two other neighboring frequencies of 86 and 690 GHz at which the modeled source could be observed.
We conclude that the results presented in the remainder of this work do not strongly depend on the chosen GRMHD grid resolution. Around the current level of resolution, which is typically used for building EHT GRMHD model libraries, all the modeled quantities are rather settled, and the small variations in observables are due to different realizations of turbulence in individual runs [^1] Turbulence is always initialized by a random perturbation of the gas internal energy.. We note that we cannot exclude that our models are located at the resolution plateau and that increasing resolution 10 times in each direction may result in the simulation better resolving more physical effects (e.g., better resolution of shocks and reconnection layers may lead to different electron heating in different regions of the flow) that could lead to quantitative and qualitative changes in the emission properties.
@@ 3 Results
3.1 Turbulent heating vs. ($R~{}(\beta)$) model
The parameters of all the fiducial runs are summarized in Table 1. Figure 3 shows an example of snapshots from our fiducial two-temperature MAD $a_{*}=0.9375$ simulation. The fiducial lightcurves produced by the GRMHD simulations with different parameters are presented in Appendix A and here we discuss distributions of quantities calculated based on these lightcurves.
Figure 4 displays the correlations of the observables in models R1 - 160 and K for the default viewing angle $i=160^{\circ}$, combined for black hole spins $a_{*}=0,0.5,0.9375$. In the $M_{3}$ panels, the K models are moderately correlated with the R1 - 160 models, and the K models are generally less variable compared to models R1 - 160. The spectral slope $\alpha_{I}$ and CP in model K are highly correlated with those of model R10. This is not the case for RM and LP, for which the correlation is much weaker.
Figure 5 shows the comparison of the distributions of the observables for the models shown in Figure 4. The averages of $M_{3}$, LP, and CP (and standard deviations) in model K are weakly distinguishable from those of models R1 - 160.
It is useful to look at the properties of the maps used to synthesize the integrated observables. Figure 6 shows examples of EHT-like images at 229 GHz produced by a random snapshot of the $a_{*}=0.5$ simulation with R1, R10, R40, R160 and K electron models for the fiducial viewing angle of $i=160^{\circ}$. Following Event Horizon Telescope Collaboration et al. 2024 we characterize these images using the following metrics. Net LP and CP are defined as
@@
The image-averaged LP and CP are defined as
@@
where sums are carried out over all image pixels. The geometry of LP is described with amplitudes and phases of a complex function,
@@
is the complex polarization vector and $\rho,\phi$ are the polar coordinates in the image plane (see Palumbo et al. 2020 or Event Horizon Telescope Collaboration et al. 2024 for details). All these metrics are shown in the example snapshot in Figure 6. Here, it is more evident that the resolved images of model K are most similar to model R10 in both LP and CP. Model R1 differs from K in CP maps. Models R40 - 160 already have a distinct appearance compared to model K in both LP and CP, as also evident from the comparison of integrated CP. When scored against existing or future (total intensity or polarimetric) EHT data, model K may favor similar accretion flow and geometrical parameters as model R10.
It is important to show that the $T_{\rm p}/T_{\rm e}$ in the K models does not precisely follow the $R~{}(\beta)$ model R10, regardless of the fact that $Q^{+}_{\rm p}/Q^{+}_{\rm e}$ is strongly dependent on $\beta$. An example of $T_{\rm p}/T_{\rm e}$ maps for a single snapshot of the MAD model with high spin is shown in Figure 7 (top and middle panels). The discrepancy between $T_{\rm p}/T_{\rm e}$ in models R10 and K indicates that there is no simple relationship between $Q^{+}_{\rm p}/Q^{+}_{\rm e}$ and $T_{\rm p}/T_{\rm e}$ as already shown by Ressler et al. (2015) for non-MAD models. Hence, the exact electron temperature consistency is not the main reason behind the similarity of K and R10 models. The difference in $T_{\rm p}/T_{\rm e}$ between the models is the smallest close to the black hole, and it increases with radius. The emitting regions close to the black hole have on average $T_{\rm p}/T_{\rm e}$ =10. The R10 and K models therefore share the scaling unit ${\mathcal{M}}$ to reproduce the same flux, meaning that $\rho$ and $B$ in the synchrotron radiative transfer coefficients are identical, leading to similar images. The differences between the model’s emissions reflect differences in $\Theta_{e}$ (shown in the bottom panels of Figure 7) and seem small from the observational point of view. Due to the Faraday rotation effects, LP and RM are the most sensitive among our observables to electron temperatures at small and larger radii. This explains the weaker correlation of LP and RM between the R10 and K models visible in Figure 4.
@@ 3.2.1 Effects of radiative cooling
We perform a GRRMHD simulation of the MAD $a_{*}=0$ model to check the impact of radiative cooling on electron temperatures and emission properties. The radiative processes included in the GRRMHD simulation include synchrotron emission, self-absorption, Compton scatterings, and bremsstrahlung. To reduce the large computational cost of models with activated multifrequency radiative transfer, the interactions between the photon field and the gas are allowed only within $100$M. Any radiative transfer effects outside of this sphere are negligible. When producing emission, the code takes into account the fact that the combined four-momentum of gas and radiation has to be conserved. Emission/absorption/scattering can therefore exert force on the gas and change the electron thermodynamics (radiative cooling or heating). In the models considered here, the average optical thickness of the gas is $\tau\ll 1$ and luminosities are strongly sub-Eddington, $L\ll L_{\rm Edd}$; hence, it is expected that in models with activated radiative transfer, mainly electron thermodynamics is affected and mainly due to synchrotron cooling. Other effects are weak for the accretion rates considered. In Appendix B, we present quality factors for the GRRMHD simulation to demonstrate that the interactions between gas and radiation are resolved by an adequate number of Monte Carlo samples.
@@
Figure 8 displays volume-averaged radial profiles of density, magnetic field, $T_{\rm p}/T_{\rm e}$ (or $R$), and $\Theta_{e}$ in the GRMHD and GRRMHD model with the K heating mechanism and ${\mathcal{M}}=5\times 10^{11}$ at different time moments, earlier and later in the evolution. All quantities are additionally weighed by the $\rho B$ factor to show quantities averaged over the regions where most of the synchrotron emission comes from (the $\rho B$ factor is a major factor in synchrotron emissivity; see, e.g. Pandya et al. 2016 and their Eq. 24 or Eq. 29). There is a small difference in the radial profiles $\Theta_{e}$ and $R$ in models with and without radiative effects, and the difference is due to the different realization of turbulence in each model. Unsurprisingly, at later times in both models, $R\sim 10$ for $r<20$M, consistent with the results in Section 3.1. Figure 9 shows that there is no difference in radiative characteristics between the GRMHD (K) and GRRMHD (Kcool) models based on $15,000-30,000$M light curves. The total 230 GHz flux of model K averaged over the above time period is 2.99 and 2.93 Jy in models with and without radiative effects, respectively. This suggests that the radiating models simply do not cool fast enough to significantly affect the electron temperatures. This is expected considering very small accretion rates onto Sgr A*.
@@
Since the test model with cooling has the highest ${\mathcal{M}}$ among all K models considered in this work (see Table 1), all other K models shown in the paper are expected to cool down even slower than the test case. Radiative cooling effects in MADs with turbulent heating may be neglected when modeling Sgr A* millimeter emission.
3.2.2 Effects of nonthermal electrons
@@
Figure 10 shows the comparison of the observational characteristics of thermal and nonthermal K fiducial runs. Non-thermal models are slightly less variable compared to the thermal models. The spectral slope $\alpha_{I}$ and the RM and the LP of both thermal and nonthermal models are strongly correlated (the correlation is much stronger compared to the $R~{}(\beta)$ models). It is evident that the total intensity and LP are not sensitive to details of the distribution function for the considered part of the electromagnetic spectrum. CP, on the other hand, is different in both models; hence it may be used to discriminate between distribution functions. We note that CP is usually a more permissive observational constraint compared to LP (see, e.g., Event Horizon Telescope Collaboration et al. 2024). Overall, with the exception of CP, all radiative properties of our simple nonthermal models are tightly correlated with their purely thermal counterparts.
@@
Having checked how the results depend on numerical resolution effects and the two most relevant physical effects, we can now compare the thermal two-temperature models directly to observations of plasma around Sgr A*.
Figure 11 shows the comparison of the prograde models K, at various viewing angles, to Sgr A* observables detected by the Atacama Large Millimeter/submillimeter Array (ALMA) at 86 and 213 - 229 GHz in 2017, April (Wielgus et al., 2022a, b, 2024). The modulation index at both frequencies increases with black hole spin and inclination angle. In general, all models are too variable to match the observed variability (this is also the case for the R1 - 160 models found in Event Horizon Telescope Collaboration et al. 2022 and shown here in Figure 5), but the data strongly favor lower spins at lower inclination angles, for which $M_{3}$ is the smallest. All models are too optically thin when compared to observations (as is also the case for the R1 - 160 models; compare our Figure 5 with Figure 12, and see also Ricarte et al. 2023). All models also require an external Faraday screen to explain the observations of RM, but we notice that models with low viewing angles ($i=160^{\circ}$, $150^{\circ}$) best recover the RM standard deviation toward the source. LP of the $i=160^{\circ}-150^{\circ}$ models is consistent with both 229 and 86 GHz measurements. Models with $a_{*}=0$ recover CP at 229 GHz but fail to recover CP at 86 GHz. On the other hand, models with $a_{*}=0.9375$ fail to recover CP at 229 GHz, but do recover CP at 86 GHz. Only the model with $a_{*}=0.5$ does reasonably well with regard to CP at both frequencies. It is possible that a model with a black hole spin in between 0 and 0.5 would be a better fit to the ALMA CP data. Among prograde MADs, the model with $a_{*}=0.5$ is the closest to the ALMA observations overall.
Although turbulence in the retrograde models reached a quasi-stationary state within the emission zone, they are reported separately because they are integrated over shorter times compared to the prograde models. Figure 12 compares the statistical properties of the light curves produced by retrograde MADs. The two-temperature MAD with $a_{*}=-0.5$ performs better compared to MAD $a_{*}=-0.9375$ as they are less variable, but worse than MAD $a_{*}=0.5$, because prograde models better recover CP at both frequencies.
@@ 4 Discussions
In this work, we have carried out several two-temperature MAD simulations with the main purpose of investigating the impact of various numerical parameters on the emission properties of these models, in particular grid resolution, but we also investigate the impact of physical effects.
@@
Regarding physical parameters, radiative cooling can be safely omitted in Sgr A* two-temperature MAD models with turbulent heating, as radiative effects do not significantly affect the electron temperature evolution. These effects may be important when modeling M87, which is a brighter source. We have also carried out the first exploratory survey of models in which turbulent heating is dissipated into a nonthermal electron distribution function. Our nonthermal models display very similar characteristics compared to the thermal models. However, including the physics of nonthermal electrons in the emission models slightly decreases the model variability and alters the CP compared to purely thermal models. Other, also spatially variable parameters of the $\kappa$ distribution function could be considered and change our result (e.g., Davelaar et al., 2019).
@@
Long-duration two-temperature simulations indicate that the thermal models of K are only roughly approximated with the R10 models. Using total intensity observables, Mizuno et al. (2021) found that the best match to turbulent heating models are R5 models. Given that their electron temperatures are defined differently (compare their Eq. 8 with our Eq. 4) and are a factor of $\sim 2$ smaller than ours, the best match with R10 found here is consistent with the previous finding.
@@
In Section 3.3 we show that none of our long-duration two-temperature K simulations recover the general properties of light curves of Sgr A* observed by ALMA ideally, but the ALMA data favor models with lower viewing angles $i=150^{\circ}-160^{\circ}$ for which the variabilities of Stokes ${\mathcal{I}}$ and RM are the smallest. All models have great difficulty recovering the observed model spectral index, in excellent agreement with the previous two-temperature study by Dexter et al. 2020a. All models recover the observed net LP but require the introduction of an additional external Faraday screen to recover the observed RM. Only the $a_{*}=0.5$ model recovers Sgr A* net CP at two observed frequencies simultaneously, and it is possible that a spin between 0 and 0.5 would produce even better results. The preferred moderate/low spin value is consistent with spin estimates based on two-temperature models of Dexter et al. 2020a, although for different than the turbulent heating scenario, different initial conditions, and different observational datasets. Here we also study the evolution of two-temperature MADs for low and high spins for a significantly longer time compared to Dexter et al. 2020a.
@@
It is possible that ALMA observes varying compact emission together with less varying extended emission that is not captured in our simulations (C.F. Gammie 2025, private communication); hence, it is preferable to compare models with EHT images (or best with the future EHT movies) rather than integrated quantities. Compared to EHT data, the K models are expected to favor a parameter space similar to the R10 models. Polarimetric EHT data of Sgr A* have not yet ruled out all MAD R10 models. In fact, 10 out of 20 MAD models passing are R10 models (see Fig. 8 in Event Horizon Telescope Collaboration et al. 2024). Based on the similarity between the appearance of R10 and K, we argue that the turbulent electron heating scenario could still be consistent with the observations of Sgr A*. It is interesting to note that in Event Horizon Telescope Collaboration et al. 2024, MAD model R10 is favored for a set of intermediate viewing angles ($i=150^{\circ}-110^{\circ}$) and positive and negative spin values. K models could narrow the range of these two geometric parameters.
@@
Based on joint total intensity and polarimetric EHT data, the current best-bet MAD model for Sgr A* is R160, not R10 (Event Horizon Telescope Collaboration et al., 2024). We should discuss the present results in the context of other models for $Q^{+}_{\rm p}/Q^{+}_{\rm e}$. As shown by Dexter et al. (2020a) assuming the Howes (2010) turbulent heating model for $Q^{+}_{\rm p}/Q^{+}_{\rm e}$, there is a negligible difference in the MAD electron temperatures compared to the Kawazura et al. (2019) model. The reconnection $Q^{+}_{\rm p}/Q^{+}_{\rm e}$ model of Rowan et al. 2017, on the other hand, may result in slightly lower (but also around R10) electron temperatures with weaker plasma-$\beta$ dependency (Dexter et al., 2020a; Mizuno et al., 2021). Lower electron temperatures could make the MAD models optically thicker and less varying (Chan et al., 2024), thus pushing them toward the right direction, but this remains to be thoroughly examined. Weaker dependence on $\beta$ in models with reconnection heating can result in more disklike emission of MADs than disk/jetlike emission (Chael et al., 2018), making it more difficult to explain the nearly flat radio spectrum of Sgr A* (Mościbrodzka & Falcke, 2013). The radio spectral slope dependency on $Q^{+}_{\rm p}/Q^{+}_{\rm e}$ should be carefully checked with high-resolution (jet-resolving) two-temperature MAD models. However, whichever $Q^{+}_{\rm p}/Q^{+}_{\rm e}$ scenario above (or a combination of scenarios) is selected, the two-temperature MAD models will have difficulty in naturally developing R160-like electron temperatures near the event horizon. Assuming that MAD is present in the Sgr A* system, this may suggest that $R~{}(\beta)$ models are inappropriate for MADs (we briefly discuss prospects on how to resolve this issue in Appendix C). Alternatively, comparisons of instantaneous model snapshots with time-averaged EHT images synthesized from the time-varying EHT data sets so far (see Event Horizon Telescope Collaboration et al. 2022, 2024) may be introducing some bias in the inference of the electron temperature parameter.
@@
Our simulations are run with constant adiabatic index $\gamma_{\rm ad}=13/9$, which is not a self-consistent assumption in regions where electrons are mildly or subrelativistic. Recent simulations of Narayan et al. (2022) suggest that the MAD dynamics is not affected by the choice of $\gamma_{\rm ad}$. It is unclear how different, or varying, $\gamma_{\rm ad}$ alters the discussed total intensity and polarimetric radiative properties of two-temperature MADs. This topic should be investigated in the future.
The recent manuscript by Salas et al. (2024), which was posted after this paper was submitted, models Sgr A* variability in total intensity with two-temperature $a_{*}=0.94$ MAD models combining reconnection heating, varying $\gamma_{\rm ad}$ index, and radiative cooling. Although we agree with their statement that the physical two-temperature models are typically less variable compared to the $R~{}(\beta)$ models (via $M_{3}$ characteristics; see our Figure 5, second top panel), further detailed comparisons of our results to Salas et al. (2024) are difficult due to different assumptions in electron heating model, equation of state, and radiative cooling methodology. Interestingly, Salas et al. (2024) models, as ours, still do not produce variability fully consistent with the observational data of Wielgus et al. (2022a).
@@
The ideal GRMHD models considered here do not include many other potentially important effects. Additional physics, such as resistivity (e.g., Vos et al., 2024), viscosity, plasma composition (Emami et al., 2021; Wong & Gammie, 2022), anisotropic pressures (e.g., Foucart et al., 2017), or anisotropic electron distribution functions (e.g., Galishnikova et al., 2023), may all impact the radiative characteristics of MADs. The study of the influence of all these effects on observed quantities goes beyond the scope of the present work. The current simulation set constitutes a reference point for future investigations.
@@
The authors thank Maciek Wielgus for providing ALMA data and comments on the paper. We also thank Andrew Chael, Charles Gammie, and Ben Ryan for many of their comments. We gratefully acknowledge the HPC RIVR consortium (www.hpc-rivr.si) and EuroHPC JU (eurohpc-ju.europa.eu) for funding this research by providing the computing resources of the HPC system Vega at the Institute of Information Science (www.izum.si). Software ebhlight (Ryan et al., 2015, 2017, 2018), ipole (Mościbrodzka & Gammie, 2018), python (Oliphant, 2007), matplotlib (Hunter, 2007).